A man is presently 4 times the age of his son. He learned that 3 years ago he was 7 times the age of his son. How old are they now?
Answer by alicealc(293)

let:
Man's age = m years old
Son's age = s years old

m = 4*s

m - 3 = 7*(s - 3)
4s - 3 = 7s - 21
21 - 3 = 7s - 4s
18 = 3s
s = 18/3 s = 6
m = 4s = 4*6 = 24

So, now, the man is 24 years old, and his son is 6 years old